# A Multiobjective Mathematical Model for Optimal Irrigation Water Allocation

### v1

@@
Abstract Water resource management is vital for sustainable irrigation and food security under growing climatic and economic pressures. In this study, we develop two single-objective optimization models for irrigation planning, one that maximizes net benefit and the other that minimizes environmental flow deficiency, and benchmark them against previous works. We then extend the analysis to a multiobjective linear programming formulation solved through different approaches to evaluate trade-offs. Numerical experiments on the Muhuri Irrigation Project reveal three outcomes: (i) a complete scenario view with profits ranging from $0.2\times 10^{9}$ to $1.497\times 10^{9}$ and environmental flow deficits between 0 and 1200 GL, where the 1200 GL represents the theoretical annual maximum under a 100 GL uniform monthly target; (ii) explicit trade-offs showing higher profits correspond to greater ecological shortfalls; and (iii) an integration-based approach producing nearly 1000 Pareto-optimal solutions within seconds, greatly outperforming earlier studies.
Keywords Irrigation Water Optimization, Environmental Flows, Crop Allocation, Pareto front, Muhuri Irrigation Project.
@@ 1 Introduction
Although approximately 70 percent of the Earth’s surface is covered by water, only 2.5 percent constitutes fresh water, and less than 1 percent of that is accessible and usable for industrial, agricultural, and domestic purposes [18]. Specifically, only 0.0002 percent of water is available as surface water runoff, 0.62 percent exists as groundwater, and just 0.001 percent is stored in the atmosphere [16]. This severe imbalance between abundance and usability makes water scarcity one of the most critical global challenges of the present era. A sustainable future depends on effective resource management. Water resource management, therefore, requires a comprehensive and well-integrated approach. This approach should include the development of appropriate infrastructure, equitable distribution mechanisms, and the preservation of environmental flows necessary for rivers and ecosystems. Achieving these objectives is not possible without the support of robust institutions, advanced technologies, and sincere cooperation across all levels of society. Climate change, urban growth, and expanding agriculture influence the raising of water demand. Regarding this, managing water wisely has become more important than ever. Different regions face different challenges, so strategies must be tailored to local climates. For example, arid countries like Iran, Iraq, and Saudi Arabia must find ways to deal with severe water shortages and their heavy reliance on groundwater [21, 32]. In semi-arid regions like South Africa and Spain, irrigation systems need to be more flexible and drought-resistant [6, 10]. In contrast, tropical monsoon countries such as Bangladesh face the dual challenge of managing both excess rainfalls during monsoons and acute shortages in dry winter periods[28, 26].
@@
A significant initiative aimed at improving water management in southeastern Bangladesh is the Muhuri Irrigation Project (MIP). This project was developed with a dam-regulator system to support dry-season irrigation and to prevent saline water from intruding inland from the Bay of Bengal. However, it has consistently underperformed due to system inefficiencies [25].Despite its strategic national importance, the MIP has received limited attention in academic research. Determining the optimal allocation of water for each crop, along with identifying suitable crop patterns under relevant constraints, remains a pressing research challenge.
Furthermore, in the previous study by Ullah and Nehring[28] reported a maximum Environmental Flow Deficiency (EFD) of 35 GL under a uniform monthly environmental flow target of 100 GL. Since EFD is defined as the nonnegative difference between the target environmental flow and the actual environmental flow, the theoretical maximum deficit occurs when no water is allocated to the environment. In such a case, the deficit would be equal to the full annual target of 100 GL × 12 months = 1,200 GL. Therefore, the reported value of 35 GL represents less than 3 percent of the possible range of deficits. This suggests that their model captured only a narrow portion of the feasible solution space and failed to approximate the full Pareto frontier of trade-offs between agricultural benefits and ecological requirements. As a result, the ecological shortfalls are understated, and the diversity of potential optimal solutions remains unexplored. The analysis is being extended to solve two single-objective problems individually: maximizing economic benefits and minimizing environmental flow deficits, which were not examined in previous studies. By employing alternative deterministic optimization approaches, including interior-point and sequential quadratic programming across different algorithms, we seek to generate accurate and reproducible outcomes that correct unrealistic allocations and reduce computational burden. These contributions not only enhance the validity of ecological and economic trade-offs but also extend the practical applicability of multiobjective water allocation models in resource-constrained irrigation projects.
@@
    Develop two single-objective mathematical models under a common set of constraints, including pumping balance, minimum area requirements, and total land availability. The objectives are formulated and solved individually: (a) maximization of economic benefit, and (b) minimization of environmental flow deficiency.
@@
    Formulate single-objective mathematical models and obtain optimal solutions using interior-point and sequential quadratic programming across different algorithms.
@@
To address these pressing challenges, this study proposes a comprehensive optimization framework. The framework (i) evaluates and contrasts multiple optimization algorithms, and (ii) develops a robust decision-support system that closely reflects the real-world conditions faced in the field. The findings are expected to assist policymakers, irrigation planners, and water resource managers in making informed, data-driven, and climate-resilient decisions that enhance agricultural water efficiency and sustainability.
@@ 2 Decision and Managerial Insights
The proposed multiobjective irrigation optimization model is well-suited for practical use in real-world irrigation planning the Muhuri Irrigation Project and similar initiatives, not just in Bangladesh but around the world. Combining seasonal water availability, the model offers a direct translation to field-level operations. It will assist the irrigation authorities to plan for different seasons, and allocate resources efficiently. The proposed irrigation optimization model serves a range of stakeholders. Farmers can use it to choose crop plans that balance income and water use, while irrigation managers and engineers can better schedule and regulate water flows. It also helps policymakers craft sustainable water-sharing strategies, and gives researchers a flexible tool to apply in other regions or for studying climate resilience. There are several practical benefits that can be gained by this model. First, the local community would be able to improve the crop productivity, and promote the equal sharing of the water resources. Economically, the model enhances profitability by making decision of minimum area of land and amount of water should be used for each crop, while concerning the ground water usage. Ecologically, it contributes to healthier river ecosystems by helping to reduce Environmental Flow Deficiency (EFD). For researchers, the model provides a replicable framework that can be adapted to other semi-arid regions facing similar challenges. This proposed model compared established multiobjective methods, e.g., gamultiobjective, pareto search, and scalarization techniques, with interior-point,sequential quadratic programming (SQP) which provides faster and more accurate optimization results. Unlike many traditional approaches, it takes into account several constraints in practical irrigation planning. By using both surface water and groundwater in a balanced way, the model helps reduce the cost of pumping water. It also helps protect crops from being damaged by overuse or underuse of water. In the end, this saves money by using land and farming inputs more wisely. The model is built to be practical and easy to use in real-life settings. It works well with common tools like AMPL, MATLAB, or Python. The required data are easy to collect—such as crop water needs, and seasonal water flow. Since it is designed in a simple scalarized form, it can also fit smoothly into the systems already used by water management authorities. It encourages fair sharing of water among farmers, which reduces conflicts and supports stronger, more resilient rural communities.
@@ 3 Literature Review
Efficient irrigation water management has become increasingly critical in balancing agricultural productivity with environmental sustainability. Modern research focuses not only on maximizing net benefits for farmers but also on ensuring long-term ecological health by minimizing environmental flow deficiency (EFD). This dual-objective perspective guides the evolution of optimization models in water resource planning. One of the earliest comprehensive efforts in multiobjective irrigation modeling was proposed by Xevi and Khan[31], who developed a model that simultaneously maximized farm income and minimized both groundwater extraction and variable costs. This dual-objective framework laid the groundwork for many subsequent models. Following a similar path, Lalehzari et al.[12] investigated optimal crop planning under water scarcity and emphasized how water use efficiency (WUE) could be improved through seasonal adjustments and pricing strategies. Their findings showed that vegetable crops consistently delivered the highest WUE. In the Bangladeshi context, Ullah and Nehring [28] applied a multiobjective approach to the Muhuri Irrigation Project (MIP), aiming to enhance economic returns while maintaining environmental sustainability. Although the model successfully produced 34 non-dominated solutions, it required extensive computation time. This gap reveals a need for more practically grounded frameworks in the context of Bangladeshi irrigation systems. In recent years, the inclusion of environmental objectives has become more prominent. Zeinali et al. [32] developed a model that balances surface and groundwater use across dry and wet periods, which led to reduced groundwater level decline—a key component of EFD minimization. Similarly, Hoque et al. [9] emphasized the importance of preserving environmental flows, particularly in dam-influenced areas, underscoring the need to harmonize dam operations with ecological flows. Adding practical constraints, Wardlaw and Bhaktikul [30] incorporated field-level soil moisture and canal capacity into their model, making it suitable for real-world implementation. Sadati et al. [24] examined both full and deficit irrigation strategies under four probabilistic weather conditions. They found that deficit irrigation consistently improved net farm income and total cropped area—contributing directly to benefit maximization while maintaining moderate water usage. Nguyen et al. [22] introduced a decision framework to reduce the computational search space in water allocation models. Their benchmark case study showed that this approach could improve solution efficiency, especially in complex, resource-intensive scenarios. Mohammadrezapour[19] evaluated different cropping options and reported that wheat provided the highest return-to-water ratio due to its low water requirement and high yield. At the same time, crops like sugar beet, which require high water input, saw significant decline in all modeled scenarios. His findings highlighted how crop selection impacts both water use and farm income. Similarly, Torabi [27] presented an integrated water operation model that improved water reliability, reduced groundwater drawdown, and minimized aquifer overuse, ensuring both short-term productivity and long-term resource sustainability. On a national scale, Musa [21] developed a model for Saudi Arabia that forecasts a shift from groundwater and surface water to treated wastewater and desalinated water between 2020 and 2050. The model projects water cost to rise from 13 Billion AUD to 22 Billion AUD, highlighting the importance of sustainable water reuse and quality management in long-term policy planning. In a more field-level approach, Bhatia and Rana [1] emphasized optimal crop allocation, resource efficiency, and sustainable practices like crop rotation to maintain soil health and ensure economic viability for farmers. While not entirely aligned with dual-objective modeling, some studies still offer useful insights. Ikudayisi [10], working in Baghdad’s Alrustumia district, analyzed the economic return from different types of treated wastewater. Though focused more on net benefit than environmental flow, the study demonstrated how wastewater reuse can support high-value crops. Marzban [15] proposed a flexible model for handling multiple goals in water-scarce regions but did not include explicit EFD constraints. Montazar [20] applied a model in Iran’s Qazvin Irrigation Command Area to evaluate cropping scenarios involving surface and groundwater use. While relevant in hydrological terms, the model lacked a strong focus on ecological sustainability. Some authors offer broader perspectives or critiques rather than direct model contributions. Lewis and Randall [13] highlighted limitations in earlier models, such as inadequate linkage between groundwater use and crop-specific water needs. Mainuddin [14] noted that Boro rice accounts for a major portion of groundwater use in Bangladesh, but did not incorporate optimization-based planning. Boah [2] pointed to the need for integrating water quality concerns into optimization frameworks. Kovalchuk [11] focused on managing river salinization through flushing techniques, important for pollution control, though outside the scope of crop-based irrigation allocation. Given the limitations in existing models, especially in the context of the Muhuri Irrigation Project, this research proposes efficient multiobjective methods such as ’gamultiobj’, ’paretosearch’, and ’scalarization’ approaches. The model integrates economic returns, seasonal water availability, and crop-specific water requirements. This approach seeks to deliver practical, efficient, and environmentally sound irrigation planning—bridging the gap between theoretical models and field-level applications in Bangladesh.
3.1 Problem Statement
The Muhuri Irrigation Project (MIP) is located in southeastern Bangladesh, covering parts of Feni, Chattogram, and Noakhali districts. The idea behind MIP took shape in the late 1980s, when the government looked for ways to increase crop output in the region. They used surface water from the Muhuri River, sending it through a network of canals and small field channels to reach the farms. But during the dry months, when supply drops, that’s when problems start. To deal with the shortage, many farmers start using groundwater as a backup option during those dry months. The project spans roughly 23076 hectares of irrigable area out of 40080 ha [29, Figure A.3], where farmers grow a variety of crops year-round, Boro rice, wheat, maize, and vegetables being the most common. Farmers prefer Rabi crops because these require less water than Kharif crops; however, in the dry winter season, production plummeted due to only 2 percent rainfall [28]. Seasonal variability in water supply and pressure to meet production targets frequently lead to inequities, with farmers at the tail end of canals receiving less water than those near the source. Ecologically, reduced river inflow during irrigation creates Environmental Flow Deficiency (EFD), threatening aquatic ecosystems, while excessive groundwater use depletes aquifers and raises energy costs. Given these realities, there is a pressing need for an updated planning model that integrates both economic objectives and ecological considerations. A robust decision-support system is essential to guide resource allocation and ensure the long-term sustainability of the Muhuri Irrigation Project.

### v2

@@
Abstract Sustainable irrigation and food security increasingly depend on efficient water resource management in the face of growing climatic and economic constraints. In this study, we develop two single objective optimization models for irrigation planning, one that maximizes net benefit and the other that minimizes environmental flow deficiency, and compare their performance with established models reported in previous studies. We then extend the analysis to a multiobjective programming formulation solved through scalarization and genetic approaches to evaluate trade-offs. Numerical experiments on the Muhuri Irrigation Project reveal three outcomes: (i) a complete scenario view with profits ranging from $\mathdollar 0.2\times 10^{9}$ to $\mathdollar 1.497\times 10^{9}$ and environmental flow deficits between 0 and 1200 GL, where the 1200 GL represents the theoretical annual maximum under a 100 GL uniform monthly target; (ii) explicit trade-offs showing higher profits correspond to greater ecological shortfalls; and (iii) an integration based approach producing nearly 1000 Pareto optimal solutions within seconds, greatly outperforming earlier studies.
Keywords Irrigation Water Allocation, Environmental Flows, Crop Allocation, Multiobjective Optimization, Muhuri Irrigation Project.
@@ 1 Introduction
Irrigated agriculture remains one of the largest consumers of freshwater worldwide, particularly in regions where seasonal variability and competing demands place sustained pressure on limited water resources. Although approximately 70% of the Earth’s surface is covered by water, only 2.5% constitutes fresh water, and less than 1% of that amount is accessible for industrial, agricultural, and domestic use [17]. Of this total, only 0.0002% is available as surface water runoff, while 0.62% exists as groundwater, and just 0.001% is stored in the atmosphere [15]. This imbalance between water availability and practical usability has made effective irrigation water allocation a critical challenge for ensuring food security, sustaining rural livelihoods, and protecting aquatic ecosystems. In many agricultural systems, irrigation planning must reconcile competing objectives: maximizing crop production while maintaining sufficient river flows to sustain ecological processes. These challenges are intensified by climate change, population growth, and expanding cultivation, which together increase water demand while reducing the reliability of surface water supplies. As a result, irrigation systems increasingly rely on conjunctive use strategies, where surface water is supplemented by groundwater during dry seasons. Under such conditions, allocation decisions are no longer purely operational but become strategic planning problems with long term economic, environmental, and social consequences.
Environmental flow, or e-flow, refers to the quantity, timing, and quality of water required to sustain the ecological integrity and functioning of a river system. Rather than being a fixed value, environmental flow varies seasonally in accordance with the river’s natural flow regime, supporting key processes such as fish migration, floodplain inundation, wetland regeneration, and groundwater recharge. Excessive water abstraction for agriculture or industry disrupts this natural rhythm, leading to ecosystem degradation, reduced recharge, and long term environmental and social costs. Maintaining minimum river flows is therefore not only an ecological requirement but also an economic and societal necessity. Any shortfall below the required environmental flow level is defined as Environmental Flow Deficiency (EFD), the magnitude of which depends critically on how realistically and flexibly environmental flow targets are specified within planning frameworks.
These challenges are particularly evident in monsoon dominated agricultural systems. Countries such as Bangladesh experience excessive rainfall during monsoon months alongside acute water shortages in the dry winter season. Irrigation systems in these regions must therefore manage both surplus and scarcity while supporting crop production across highly contrasting hydrological conditions. Allocation decisions directly influence cropping patterns, cultivated area, seasonal yields, and farmer livelihoods, making robust planning tools essential for agricultural sustainability [24, 26].
@@
A major initiative to improve water management in southeastern Bangladesh is the Muhuri Irrigation Project (MIP). The project was developed with a dam regulator system to support dry season irrigation while preventing the intrusion of saline water from the Bay of Bengal into inland areas. However, it has consistently underperformed due to system inefficiencies[26]. Effective irrigation systems depend on sound water resource management, and in this context, the MIP was designed to utilize river water for agricultural production through planned and controlled distribution. Over the past several decades, the Bangladesh Water Development Board (BWDB) has implemented numerous water resource development and management programs, among which the Muhuri Irrigation Project represents a targeted effort to enhance agricultural productivity through improved irrigation management. Despite its strategic importance at the national level, the MIP has received limited attention in academic research. Consequently, determining the optimal allocation of water among crops and identifying suitable cropping patterns under relevant constraints remains an important and unresolved research challenge.
A key limitation of existing allocation studies for the MIP lies in their restricted representation of decision space. Previous work has evaluated irrigation performance primarily within a narrow range of environmental flow deficits. For example, Ullah and Nehring [26] reported a maximum Environmental Flow Deficiency (EFD) of 35 GL under a uniform monthly environmental flow target of 100 GL. Given that EFD is defined as the non negative shortfall between target and actual environmental flow, the theoretical maximum deficit corresponding to zero environmental release would equal the full annual target of 1,200 GL. The reported value therefore represents less than 3% of the feasible deficit range, indicating that only a limited portion of the potential economic–environmental trade-off space has been explored.
This narrow focus constrains decision relevance. When irrigation systems are pushed beyond moderate operating conditions such as during prolonged dry seasons or under policy driven environmental constraints planners lack guidance on how economic benefits evolve across the full range of possible environmental shortfalls. As a result, allocation decisions risk understating ecological stress or overestimating economic resilience, limiting their usefulness for real world agricultural planning.
To address this limitation, this study aims to improve the computational representation of trade-offs between economic and environmental objectives in irrigation allocation models, using the Muhuri Irrigation Project as a benchmark case study. Specifically, the study expands the decision space by explicitly characterizing trade-offs between net economic benefit and environmental flow deficiency across a substantially wider allocation range. By developing and comparing complementary optimization formulations under a consistent set of operational constraints, the study enables systematic evaluation of economic–environmental trade-offs under stylized seasonal water scarcity conditions.
@@
    Develop two single objective mathematical models under a common set of constraints, including pumping balance, minimum area requirements, and total land in the MIP. The objectives are formulated and solved individually: (a) maximization of economic benefit, and (b) minimization of environmental flow deficiency.
@@
    Formulate single objective mathematical models and obtain optimal solutions using interior point and sequential quadratic programming across different algorithms.
@@ 2 Decision and Managerial Insights
This study presents three mathematical optimization models for irrigation water allocation in an irrigation system that relies primarily on surface water while supplementing supply with groundwater during periods of seasonal scarcity. The models are designed to represent alternative planning priorities and to clarify trade-offs between economic performance and environmental flow protection under constrained water availability. Model 1 emphasizes the maximization of net economic benefit, highlighting allocation strategies that favor agricultural production but may increase environmental flow deficit. Model 2 prioritizes the reduction of environmental flow deficiency, supporting ecological sustainability at the expense of lower economic returns. Model 3 integrates both objectives within a multiobjective optimization framework, allowing economic and environmental outcomes to be evaluated simultaneously and supporting balanced allocation decisions.
A key contribution of this study is the expansion of the decision space beyond the limited operating ranges commonly examined in earlier planning efforts. Previous analysis on MIP largely focused on moderate levels of environmental flow deficit, restricting managerial insight to a narrow set of near optimal outcomes. In contrast, the proposed framework illustrates how net economic benefits evolve across a much wider range of environmental flow deficits within a short evaluation horizon. This extended representation enables farmers and local authorities to assess not only favorable allocation options but also the economic implications of operating under increasingly stressed conditions, which are often encountered during dry seasons.
The proposed models further enhance decision support by treating both crop land allocation and environmental flow release as endogenous decision variables rather than fixed parameters. This formulation allows the optimization process to determine optimal crop area distribution and environmental flow allocation simultaneously, subject to shared operational constraints such as total cultivable land, pumping balance, and minimum area requirements. As a result, the interaction between agricultural production and ecological protection emerges directly from the model structure rather than being imposed through predefined allocation rules.
The multiobjective formulation explicitly characterizes trade-offs between net economic benefit and environmental flow deficiency by generating a Pareto set of feasible solutions. Instead of relying on ad hoc weighting schemes, this approach allows decision makers to evaluate multiple compromise strategies and select allocation policies that align with shifting priorities across seasons, hydrological conditions, or policy objectives. By making these trade-offs transparent, the framework supports more defensible and adaptive planning decisions in contexts where economic and environmental goals compete.
From a managerial perspective, the proposed models function as a practical decision support tool for irrigation planners, water resource managers, and policymakers. The framework enables users to compare allocation strategies across the full operating range, identify zones of diminishing economic returns, and understand the consequences of pushing the system toward higher levels of environmental stress. This reduces reliance on trial and error decision making and supports quicker, more informed responses under water scarcity and institutional pressure.
The analysis is conducted entirely using secondary data previously published for the Muhuri Irrigation Project, ensuring methodological consistency and direct comparability with earlier studies. All hydrological inputs, crop parameters, and environmental flow requirements are adopted without modification. Accordingly, the contribution of this study lies in improving solution robustness, efficiency, and trade-off representation through alternative optimization formulations rather than in site specific data recalibration or operational redesign.
The scope of the study is clearly defined. While the models are calibrated for the Muhuri Irrigation Project, the underlying framework is applicable to other regulated irrigation systems that rely primarily on surface water while supplementing supply with groundwater during dry periods. At the same time, the conclusions remain grounded in local hydrological, institutional, and socioeconomic conditions, ensuring relevance without overgeneralization.
@@ 3 Literature Review
Efficient irrigation water management has become increasingly important as agricultural systems face growing pressure to balance economic productivity with environmental sustainability. Contemporary irrigation planning therefore extends beyond profit maximization to include ecological objectives, particularly the preservation of riverine health through environmental flow protection. A key metric used to represent this ecological requirement is Environmental Flow Deficiency (EFD), defined as the non negative shortfall between a target environmental flow and the actual flow released to the river. When actual releases fall below the ecological target, the deficit contributes to EFD, while exceedances do not incur penalties. This formulation provides a transparent and quantifiable way to incorporate ecological constraints into irrigation water allocation models.
One of the earliest comprehensive multiobjective optimization frameworks for irrigation planning was proposed by Xevi and Khan [29], who developed a model that simultaneously maximized farm income while minimizing groundwater extraction and variable irrigation costs. Their work established the foundation for integrating economic and hydrological objectives within a single optimization framework. Building on this approach, Lalehzari et al. [11] investigated optimal crop planning under water scarcity and demonstrated how water use efficiency could be improved through seasonal adjustments and pricing mechanisms. Their results showed that vegetable crops consistently achieved higher water use efficiency, reinforcing the importance of crop selection in water allocation decisions. These early studies collectively demonstrated the potential of multiobjective optimization in irrigation planning, although explicit ecological flow targets were not yet a central modeling focus.
Subsequent research increasingly emphasized environmental considerations in irrigation modeling. Zeinali et al. [30] developed a framework that balanced surface and groundwater use across wet and dry seasons, demonstrating that coordinated use could significantly reduce groundwater depletion, a key contributor to environmental flow preservation. Hoque et al. [8] further highlighted the importance of maintaining environmental flows in dam regulated systems, underscoring the need to harmonize dam operations with downstream ecological requirements.
At the operational level, Wardlaw and Bhaktikul [28] incorporated field scale constraints such as soil moisture dynamics and canal capacity, thereby improving the realism and applicability of irrigation optimization models. Sadati et al. [23] examined full and deficit irrigation strategies under probabilistic weather conditions and showed that deficit irrigation could improve net farm income while maintaining moderate water use. Nguyen et al. [21] contributed methodologically by proposing a framework to reduce the computational search space in large scale allocation problems, improving solution efficiency in complex systems.
Several studies examined the influence of crop selection on both economic returns and water use. Mohammadrezapour [18] reported that wheat achieved the highest return to water ratio due to its relatively low water requirement and high yield, while water intensive crops such as sugar beet performed poorly under scarcity. Torabi [25] proposed an integrated operational model that improved water reliability, reduced groundwater drawdown, and minimized aquifer overuse, emphasizing the importance of coordinated surface and groundwater management.
At a broader policy scale, Musa [20] developed a national water planning model for Saudi Arabia, projecting a long term shift toward treated wastewater and desalinated water sources. Although contextually different, the study highlighted the economic consequences of unsustainable water use. Complementing this perspective, Bhatia and Rana [1] emphasized optimal crop allocation and sustainable farming practices such as crop rotation to maintain soil health and farm viability.
Several studies provide important conceptual foundations but do not explicitly incorporate environmental flow deficiency. Ikudayisi [9] analyzed economic returns from treated wastewater reuse in Iraq, demonstrating benefits for high value crops but without explicit ecological flow constraints. Montazar [19] evaluated cropping scenarios using both surface and groundwater in Iran’s Qazvin Irrigation Command Area, though ecological sustainability was not a primary objective. Marzban [14] proposed a flexible multi goal optimization framework for water scarce regions but did not explicitly model EFD.
Lewis and Randall [12] made a key methodological contribution by proposing a multiobjective optimization framework that explicitly linked irrigation decisions, groundwater extraction, and crop specific water requirements. Their work provided the theoretical basis for subsequent applications of multiobjective techniques in regulated irrigation systems and informed later case studies in different geographical contexts.
In the Bangladeshi context, Ullah and Nehring [26] applied a multiobjective optimization framework to the Muhuri Irrigation Project (MIP), using groundwater pumping limits and target environmental flow parameters adapted from the framework of Lewis and Randall. While their study demonstrated the feasibility of applying multiobjective methods to the MIP, it required substantial computational effort and explored only a limited portion of the feasible trade-off space.
The present study builds upon this line of research while introducing several methodological advancements. Unlike earlier studies that relied primarily on multiobjective formulations, this research independently analyzes two single objective models—maximizing economic benefit and minimizing environmental flow deficiency—under a common set of operational constraints. This approach enables the identification of extreme solutions that define the boundaries of the feasible solution space. Furthermore, alternative deterministic optimization techniques, including interior point and sequential quadratic programming methods, are employed to improve computational efficiency and reproducibility. While groundwater pumping limits and environmental flow targets are consistent with earlier studies, their applicability is justified by the hydrological, institutional, and operational characteristics of the Muhuri Irrigation Project, including seasonal water scarcity, centralized management, and the need to prevent saline intrusion in a coastal river system.
Given the limitations of existing models, particularly in the context of the Muhuri Irrigation Project, this study proposes efficient multiobjective optimization implemented techniques, including ‘gamultiobj’, ‘paretosearch’, and scalarization approaches. By integrating economic returns, crop specific water requirements, seasonal water availability, and explicit environmental flow constraints within a unified optimization framework, the study advances irrigation water allocation modeling and bridges the gap between theoretical formulations and practical, field level decision making in Bangladesh’s irrigation sector.
4 Problem Statement
Despite these objectives and its socioeconomic importance, the performance of the MIP has fallen short of initial expectations. During the dry winter months, surface water availability from the Muhuri River declines sharply, while rainfall contributes only about 2% of the annual total [26]. Although farmers tend to prefer Rabi crops because of their lower water requirements compared to Kharif crops, the limited availability of surface water during this period restricts the effective use of irrigation. As a coping strategy, many farmers increasingly depend on groundwater extraction, which raises production costs and places additional pressure on already stressed aquifers.
Institutional and infrastructural limitations further exacerbate these challenges. Inadequate water distribution systems, poor drainage conditions, and unplanned cropping intensities reduce the effectiveness of available irrigation water. Moreover, the centralized operation of the canal network, combined with limited control over on farm water delivery, often results in inequitable access to water, particularly disadvantaging farmers located at the tail ends of canals. Competition among users intensifies during dry periods, increasing the likelihood of inefficient and uneven water use. From an ecological perspective, reduced river inflows caused by irrigation withdrawals contribute to Environmental Flow Deficiency (EFD), threatening aquatic ecosystems and local fisheries that depend on sufficient freshwater availability. At the same time, excessive reliance on groundwater undermines long term sustainability by accelerating aquifer depletion and increasing energy consumption.
These interconnected socioeconomic, institutional, and environmental challenges highlight the limitations of existing operational practices within the Muhuri Irrigation Project. They underscore the need for an updated and integrated planning framework that explicitly accounts for both economic objectives and ecological constraints. A robust decision support system capable of guiding irrigation water allocation under conditions of seasonal scarcity is therefore essential to improve equity, provide a comprehensive representation of Environmental Flow Deficiency, support informed decision making by local authorities, enhance agricultural productivity, and ensure the long term sustainability of the Muhuri Irrigation Project.